# 3D Gaussian Splatting: Survey, Technologies, Challenges, and Opportunities

### v1

@@ 6.3 Splatting
The role of Splatting is to efficiently transform 3D Gaussian data into high-quality 2D images, ensuring smooth, continuous projections and significantly improving rendering efficiency. As a core technology in traditional computer graphics, there are also efforts aimed at improving it from the perspectives of efficiency and performance [137].
Figure 7: The Modified Splatting Strategy [190].
Introducing the ADOP [12] in 3DGS for real-time rendering, TRIPS [191] utilizes a screen-space image pyramid for point rasterization and employs trilinear write for rendering large points. This approach, combined with front-to-back alpha blending and a lightweight neural network for detail reconstruction, ensures crisp, complete, and alias-free images.
The work [190] identifies limitations in prior methods [15] that use local affine approximations during the projection process, resulting in errors detrimental to rendering quality. By analyzing the residual errors from the first-order Taylor expansion, they establish a correlation between these errors and the Gaussian mean position. Based on this analysis, the unified projection plane is replaced with tangent planes, mitigating the projection errors in the original 3DGS through the Unit Sphere Based Rasterizer, as shown in Fig. 7. A similar issue is also mentioned in 2D-GS [137], which is introduced in Sec. 5.2.
6.4 Regularization
Regularization is crucial for 3D reconstruction. However, due to the lack of direct supervision from 3D data, the original 3DGS supervises its training by comparing the rendered images with the real images. This form of supervision can lead to training instability, particularly in under-determined scenarios such as those with sparse view settings. In this section, we categorize the regularization terms into 2D and 3D regularization terms, as shown in Fig. 8. The 3D regularization primarily provides constraints for 3DGS directly, whereas the 2D regularization terms impose constraints on the rendered images, thereby influencing the optimization of attributes.
@@ 6.4 Regularization
6.4.1 3D Regularization
The 3D regularization has garnered significant attention due to its intuitive constraint capabilities. These efforts can be categorized based on their targeted objectives into individual Gaussian primitive, local, and global regularization.
Individual Gaussian Primitive Regularization: This regularization primarily aims to improve certain attributes of Gaussian primitives [126, 32, 139, 78], as mentioned in Sec. 6.2.
Local Regularization: Owing to the explicit representation of 3DGS, it is meaningful to impose constraints on Gaussian primitives within local regions. Such constraints can ensure the continuity of Gaussian primitives in the local space. Physics-related Regularization is often used to ensure the local rigidity of deformable targets, which includes short-term local-rigidity loss, local rotation similarity loss, and long-term local isometry loss. Short-term local rigidity implies that nearby Gaussians should move following a rigid body transformation between time steps,
@@
[135, 188, 150, 146, 103, 104, 147, 174]. Subsequently, some works have also adopted similar paradigms to constrain local rigidity [149, 111].
In addition to rigidity loss, Conservation of Momentum Regularization can also be used as a constraint for dynamic scene reconstruction. It encourages a constant velocity vector and applies a low-pass filter effect to the 3D trajectories, thereby smoothing out trajectories with sudden changes [146]. In addition, there are some Local Consistency Regularization terms that also aim to constrain the Gaussian primitives within local regions to maintain similar attributes, such as semantic attributes [176, 138, 179], position [98, 107], time [149], frame [141], normal [192] and depth [52].
Global Regularization: Unlike the local regularization within neighboring regions, global regularization aims to constrain the overall 3DGS. Physics-related Regularization introduces real-world physical characteristics to constrain the state of 3DGS, including gravity loss and contact loss, among others. Gravity loss is used to constrain the relationship between the object and the floor, while contact loss targets the relationships among multiple objects [98].
Benefiting from the explicit representation of 3DGS, depth and normal direction attributes can be directly obtained, which can serve as constraints during training, particularly for surface reconstruction tasks. Depth-Normal Regularization achieves depth-normal consistency by comparing the normal computed from the depth values with the predicted normal [160, 37, 158, 161]. This method effectively enforces constraints on both normal and depth simultaneously. Additionally, directly constraining either the normal or the depth is also feasible.Normal Regularization often adopts a self-supervised paradigm due to the lack of direct supervision signals, which can be implemented by designing pseudo-labels from Gradient [153], Shortest axis direction of Gaussian primitives [37] or SDF [159, 158]. Similarly, Depth Regularization also adopts a similar approach; however, it not only aims for accurate depth values but also seeks to ensure clear surfaces in 3DGS. Depth Distortion loss [8] aggregates Gaussian primitives along the ray,
@@
where $z$ is the intersection depth of Gaussian [137, 161]. In addition to self-supervised methods, incorporating additional pre-trained models to estimate normal [160] and depth [38, 56, 54, 193] has proven to be more effective in Normal Regularization and Depth Regularization. Based on them, derived works introduce hard depth and soft depth regularization to address the geometry degradation and achieve more complete surfaces [56]. A similar regularization term is SDF Regularization, which is also a constraint strategy for surface reconstruction. It achieves the desired surface by constraining the SDF that corresponds to 3DGS to an ideal distribution [156, 153, 159, 158, 194].
6.4.2 2D Regularization
Unlike the intuitive constraints of 3D regularization, 2D regularization is often used to address under-constrained situations where original loss functions alone are insufficient.
SDS loss: An important example is the SDS loss, as shown in Eq.3, which leverages a pre-trained 2D diffusion model to supervise 3DGS training through distillation paradigms  [75, 162]. This approach is also extended to distill pre-trained 3D diffusion models [82], multi-view diffusion models [195], image editing models [164], and video diffusion models. Among these, the introduction of 3D diffusion models [80, 196] and multi-view diffusion models [76, 77, 78, 103, 104, 107] can optimize the reconstruction of explicit geometry and multi-view consistency. Image editing models [163] can achieve controllable editing, while video diffusion models [103] can be used for dynamic temporal scene generation with video SDS. Based on this paradigm, the distillation of other modal images also holds potential, as it can provide more constraints from corresponding pre-trained diffusion models, such as RGB-Depth [74], where designers need to discuss how to construct diffusion models.
Some improvements specifically target inherent issues in SDS [89, 85]. Interval Score Matching is proposed to address issues of randomness and single-step sampling,
@@
where $s=t-\delta_{T}$ and $\delta_{T}$ denotes a small step size [85]. Introducing Negative Prompts [197] is also a method [74, 85, 196] to mitigate the impact of random noise $\epsilon$ and enhance stability, by replacing random noise with negative prompts terms $[\epsilon_{\phi}\left(\bm{x}_{t};y_{\text{neg}}\right)]$. Subsequently, LODS introduces LoRA [90] terms $[\frac{(1-w)\epsilon_{\psi}(\bm{x_{t}};\emptyset,t)}{w}-\frac{\epsilon}{w}]$ to replace traditional random noise $\epsilon$, thereby alleviating the impact of out-of-distribution [89].
Flow loss: Flow loss is a commonly used regularization term for dynamic 3DGS. Using the output of the pre-trained 2D optical flow estimation model as ground truth, it can render predicted flow by calculating the displacement of Gaussian primitives over a unit time and splatting these 3D displacements onto a 2D plane [115, 104, 114]. However, this approach has a significant gap, primarily because optical flow is a 2D plane attribute and is susceptible to noise. Selecting Gaussian primitives with correct depth in space and introducing uncertainty through KL divergence to constrain optical flow is a potentially feasible method [106].
Other loss: There are also some 2D regularization terms worth discussing. For example, constraining the differences in amplitude and phase between the rendered image and the ground truth in the frequency domain can serve as a loss function to aid training, thereby alleviating overfitting issues [33]. Introducing pseudo-labels for hypothetical viewpoints through noise perturbation can assist training in sparse-view settings [54]. In large-scale scene mapping, constraining the changes in attributes before and after optimization can prevent catastrophic forgetting on 3DGS [129]. Additionally, bundle adjustment is usually an important constraint in pose estimation problems [118, 119, 130].
It should be noted that, whether 2D or 3D regularization is used, overall optimization is sometimes suboptimal due to the large number of primitives in 3DGS. Some primitives often have an uncontrollable impact on the results. Therefore, it is necessary to guide the optimization by selecting important Gaussian primitives using methods such as visibility [122, 120, 134, 131].
Discussion: For a specific task, various constraints can often be introduced, including 2D and 3D regularization terms, where many regularization terms are plug-and-play and can directly enhance performance.
6.5 Training Strategy
@@
6.5.1 Multi-stage Training Strategy
Multi-stage training strategy is a common training paradigm, often involving coarse-to-fine reconstruction. It is widely used for under-determined tasks, such as in AIGC, dynamic 3DGS construction, and SLAM.
[…]
@@ 6.5.1 Multi-stage Training Strategy
In multi-objective optimization tasks, multi-stage training paradigms can enhance training stability and performance. For example, the coarse-to-fine camera tracking strategy is a common approach that first obtains a coarse camera pose through a sparse set of pixels, and then further refines the estimated camera pose based on the optimized rendering results [118, 133].
Additionally, some works aim to refine the 3DGS trained in the first stage [57, 80, 179, 170, 188, 87, 51] or endow them with additional capacities, such as semantics [180, 134] and stylization [172]. There are many such training strategies, which are also effective means to maintain training stability and avoid local optima [24]. Furthermore, iterative optimization of the final result to enhance performance is also feasible [136].
6.5.2 End-to-End Training Strategy
End-to-end training strategies are often more efficient and can be applied to a wider range of downstream tasks. Some typical works are described in Fig. 9.
@@ 6.5.2 End-to-End Training Strategy
Progressive Optimization Strategy: This is a commonly used strategy that helps 3DGS prioritize learning global representations before locally optimizing details. In the frequency domain, this process can also be viewed as a progressive learning from low-frequency to high-frequency components. It is often implemented by gradually increasing the proportion of high-frequency signals [187, 33] or introducing progressively larger image/feature sizes for supervision [17, 39, 119], which can also improve efficiency [116]. In generative tasks, progressively selecting the camera pose is also an easy-to-difficult training strategy, gradually optimizing from positions close to the initial viewpoint to those further away [77, 102].
Block Optimization Strategy: This strategy is often used in large-scale scene reconstruction, which can not only improve efficiency but also alleviate the problem of catastrophic forgetting [121, 134, 133]. Similarly, it can achieve reconstruction by partitioning the scene into static backgrounds and dynamic objects [112, 113, 114, 144]. Additionally, such an approach is also applied in AIGC and Semantic Understanding, where refining the reconstruction quality of submaps enhances overall performance [87, 181].
Unlike submaps divided by spatial regions, Gaussian primitives can be categorized into different generations along with their densification process, allowing for the application of different regularization strategies to each generation, which is an effective strategy to regulate the fluidity of different generations [162]. Categorizing the initial points into those on smooth surfaces and independent points is also a feasible strategy. By designing distinct initialization and densification strategies for each category, it can achieve better geometric representation [192]. Additionally, some works aim to design keyframes (or windows) selection strategies based on inter-frame covisibility or geometric overlap ratio in temporal data and use them for reconstructions [122, 129, 118, 126, 120, 141].
Robust Optimization Strategy: Introducing noise perturbations during training is a common method to enhance the robustness of the training process [143, 92, 57]. Such perturbations can target camera poses, timestep, and images, and can be regarded as a form of data augmentation to prevent overfitting. Additionally, some training strategies aim to mitigate overfitting and catastrophic forgetting by avoiding continuous training from a single viewpoint [121, 127].
Distillation-based Strategy: To compress model parameters, some distillation strategies use the original 3DGS as the teacher model and low-dimensional SH 3DGS as the student model. By introducing more pseudo views, they aim to improve the representational performance of the low-dimensional SH [24].
@@ 6.6 Adaptive Control
6.6.1 Densification
Densification is crucial, especially for detail reconstruction. In this section, we will analyze it from the perspectives of "Where to densify" and "How to densify". Additionally, we will discuss how to avoid excessive densification.
Where to Densification: Densification techniques typically concentrate on identifying positions requiring densification, with this process being governed by gradients in the original 3DGS, which can also be extended to the reconstruction of new targets in dynamic scenes [140]. Subsequently, the regions with low opacity or silhouette and high depth rendered error, regarded as unreliable areas, have also been considered important factors for guiding densification [129, 118, 134, 148, 32, 160], which are typically used to fill holes or improve regions with 3D inconsistencies. Some works continue to focus on improvements based on gradients, where the number of pixels covered by each Gaussian in different views is considered as weights to dynamically average the gradients of these views, thereby improving the conditions for point cloud growth [198]. Additionally, the SDF value and neighbor distance are also important criteria, with locations closer to the surface and lower compactness being more prone to densification [159, 156, 80].
How to Densification: Numerous works have improved densification methods. Graph structures are utilized to explore relationships between nodes and define new Gaussians at the center of edges based on a proximity score, thereby mitigating the impact of sparse viewpoints [54]. To prevent excessive growth in the number of Gaussians, the Candidate Pool Strategy is designed to store pruned Gaussians for densification [94]. Additionally, work [199] introduces three conservation rules to ensure visual consistency and employs integral tensor equations to model densification.

### v2

@@ VI-C Splatting
The role of Splatting is to efficiently transform 3D Gaussian data into high-quality 2D images, ensuring smooth, continuous projections and significantly improving rendering efficiency. As a core technology in traditional computer graphics, there are also efforts aimed at improving it from the perspectives of efficiency and performance.
Several studies focus on enhancing splatting mechanisms. By analyzing the residual errors from the first-order Taylor expansion, the work [240] establishes a correlation between these errors and the Gaussian mean position and an unified projection plane to mitigate the projection errors through the Unit Sphere Based Rasterizer. Analytic-Splatting [33] advances pixel-center splatting through the introduction of pixel-window-based approximation, archiving anti-aliasing. To enhance 3DGS performance in complex scenarios, a GPU-accelerated ray tracing algorithm is introduced [241] for semi-transparent particle-based representations, achieving real-time rendering with support for complex effects such as secondary rays, depth of field, and distorted cameras. Additional improvements focus on optimizing splat ordering during the blending process. StopThePop [28] proposes a hierarchical per-pixel sorting strategy, which eliminates popping artifacts and ensures view-consistent real-time rendering by accurately computing the depth of Gaussians along individual rays.
VI-D Regularization
Regularization is crucial for 3D reconstruction. We categorize the regularization terms into 2D and 3D regularization, as shown in Fig. 5. The 3D regularization directly constrains 3DGS, while the 2D regularization imposes constraints on the rendered images, thus influencing attribute optimization.
@@ VI-D Regularization
VI-D1 3D Regularization
The 3D regularization has garnered significant attention due to its intuitive constraint capabilities. These efforts categorized based on their targeted objectives into individual Gaussian primitive, local, and global regularization.
Individual Gaussian Primitive Regularization: This regularization primarily aims to improve attributes in Gaussian primitives [238, 26, 191, 98, 80], as Sec. VI-B.
Local Regularization: Owing to the explicit representation of 3DGS, it is meaningful to impose constraints on Gaussian primitives within local regions. Such constraints can ensure the continuity and feasibility of Gaussian primitives in the local space. Physics-related Regularization is often used to ensure the local rigidity of deformable targets, which includes short-term local-rigidity loss, local rotation similarity loss, and long-term local isometry loss. Short-term local rigidity implies that nearby Gaussians should move following a rigid body transformation between time steps,
@@
[169, 236, 184, 177, 133, 134, 182, 187]. Subsequently, some works have also adopted similar paradigms to constrain local rigidity [183, 144, 179].
In addition to rigidity loss, Conservation of Momentum Regularization can also be used as a constraint in dynamic scene reconstruction. It encourages a constant velocity vector and applies a low-pass filter effect to the 3D trajectories, thereby smoothing out trajectories with sudden changes [177]. In addition, there are some Local Consistency Regularization terms that also aim to constrain the Gaussian primitives within local regions to maintain similar attributes, such as semantic attributes [219, 226, 221], position [126, 139], time [183], frame [172], normal [242], weight [38] and depth [62, 56].
Global Regularization: Unlike the local regularization within neighboring regions, global regularization aims to constrain the overall 3DGS. Physics-related Regularization introduces real-world physical laws to constrain the state of 3DGS, including gravity loss and contact loss, among others. Gravity loss is used to constrain the relationship between the object and the floor, while contact loss targets the relationships among multiple objects [126].
Benefiting from the explicit representation, depth and normal attributes can be directly calculated and constrained during training, particularly for surface reconstruction tasks. Depth-Normal Regularization achieves depth-normal consistency by comparing the normal computed from depth values with the predicted normal [197, 34, 194, 198]. This method enforces constraints on both normal and depth simultaneously. Additionally, directly constraining either the normal or the depth is also feasible. Normal Regularization often adopts a self-supervised paradigm due to the lack of direct supervision signals, which can be implemented by designing pseudo-labels from gradients [193], the shortest axis direction of Gaussian primitives [34], or SDF [195, 194]. Similarly, Depth Regularization adopts a similar approach; however, it not only aims for accurate depth values but also seeks to ensure clear surfaces in 3DGS. Depth Distortion loss [20] aggregates Gaussian primitives along the ray,
@@
where $z$ is the intersection depth of Gaussian [199, 198]. In addition to self-supervised methods, incorporating additional pre-trained models to estimate normal [197] and depth [35, 59, 57, 243] has proven to be more effective in Normal Regularization and Depth Regularization. Derived works introduce hard depth and soft depth regularization to address the geometry degradation and achieve more complete surfaces [59]. Similarly, SDF Regularization is also a constraint strategy for surface reconstruction. It achieves the desired surface by constraining the SDF that corresponds to 3DGS to an ideal distribution [192, 193, 195, 194, 244].
VI-D2 2D Regularization
Unlike the intuitive constraints in 3D, 2D regularization is often used to address under-constrained situations where original loss functions alone are insufficient.
SDS loss: An important example is the SDS loss, as shown in Eq.8, which uses a pre-trained 2D diffusion model to supervise 3DGS training via distillation paradigms [95, 201].
This approach extends to distill pre-trained 3D diffusion models [245], multi-view diffusion models [246], image editing models [203], and video diffusion models. Introducing 3D [100, 247] and multi-view diffusion models [96, 97, 98, 133, 134, 139, 117] enhances geometry and multi-view consistency. Image editing models [202] enable controllable edits, while video diffusion models [133] support dynamic temporal scene generation. Additionally, distillation on multi-modal images, like RGB-Depth [93], also holds potential, providing more constraints from pre-trained diffusion models.
Some improvements specifically target inherent issues in SDS [106, 104]. Interval Score Matching is proposed to address issues of randomness and single-step sampling.
@@
where $s=t-\delta_{T}$ and $\delta_{T}$ denotes a small step size [104]. Introducing Negative Prompts [248] is a method [93, 104, 247] to mitigate the impact of random noise $\epsilon$ and enhance stability by replacing random noise with negative prompts $[\epsilon_{\phi}(\bm{x}_{t};y_{\text{neg}})]$. And, LODS incorporates LoRA terms [249] $[\frac{(1-w)\epsilon_{\psi}(\bm{x_{t}};\emptyset,t)}{w}-\frac{\epsilon}{w}]$ to replace traditional random noise $\epsilon$, thereby alleviating the impact of out-of-distribution [106].
Flow loss: It is a commonly used regularization term for dynamic 3DGS and uses the output of a pre-trained 2D optical flow estimation model as ground truth. Predicted flow is rendered by calculating the displacement of Gaussian primitives over a unit time and splatting these 3D displacements onto a 2D plane [181, 134, 148, 179]. However, this approach has a significant gap since optical flow is a 2D attribute and susceptible to noise. Selecting Gaussians with correct depth and introducing uncertainty through KL divergence to constrain optical flow is a potentially feasible method [178].
Other loss: There are also some 2D regularization terms worth discussing. For example, constraining the differences in amplitude and phase between the rendered image and the ground truth in the frequency domain can serve as a loss function to aid training, thereby alleviating overfitting issues [27]. Introducing pseudo-labels for hypothetical viewpoints through noise perturbation can assist training in sparse-view settings [57]. In large-scale scene mapping, constraining the changes in attributes before and after optimization can prevent catastrophic forgetting in 3DGS [250]. Additionally, bundle adjustment is usually an important constraint in pose estimation problems [151, 152, 161].
Noted that, whether 2D or 3D regularization is used, overall updating is sometimes suboptimal due to the large number of primitives. Some primitives often have an uncontrollable impact on the results. Therefore, it is necessary to guide the optimization by selecting important primitives using methods such as visibility [160, 153, 164, 162].
Discussion: The incorporation of regularization terms serves as an effective approach to enhance the reconstruction performance of 3DGS. These regularization terms can impose constraints on various attributes of 3DGS, including geometry and spatial distribution, etc., in accordance with specific task requirements. Furthermore, in under-determined scenarios, the optimization process can be further constrained by introducing additional prior information. For a particular task, it is feasible to incorporate multiple distinct constraints simultaneously.
VI-E Training Strategy
@@
VI-E1 Multi-stage Training Strategy
Multi-stage training strategy is a common training paradigm, often involving coarse-to-fine reconstruction. It is widely used for under-determined tasks, such as AIGC, SLAM, etc..
Using different 3D representations in different training stages is a typical multi-stage training paradigm. 3DGS $\rightarrow$ Mesh (training 3DGS first, converting to Mesh, then optimizing Mesh) [95, 247, 97, 96, 202, 115] ensures geometric consistency in the generated 3D model. Additionally, generating multi-view images [140, 109, 247, 123, 204, 205, 206] in the first stage to aid reconstruction in the second stage can alleviate optimization difficulties.
@@ VI-E1 Multi-stage Training Strategy
In multi-objective optimization tasks, multi-stage training paradigms can enhance stability and performance. For example, the coarse-to-fine camera tracking strategy first obtains a coarse camera pose from a sparse pixel set, then refines it based on optimized rendering results [151, 163].
Additionally, some works aim to refine the 3DGS trained in the first stage [64, 100, 221, 210, 236, 105, 52, 117] or endow them with additional capabilities, such as semantics [222, 164] and stylization [213]. There are many such training strategies, which are also effective in maintaining training stability and avoiding local optima [15]. Furthermore, iterative optimization of the final result to enhance performance is also feasible [206, 61].
VI-E2 End-to-End Training Strategy
These strategies are often more efficient and can be applied to a wider range of downstream tasks. Some typical works are described in Fig. 6.
@@ VI-E2 End-to-End Training Strategy
Progressive Optimization Strategy: This commonly used strategy helps 3DGS prioritize learning global representations before locally optimizing details. In the frequency domain, this can be viewed as progressively learning from low-frequency to high-frequency components. It is often implemented by gradually increasing the proportion of high-frequency signals [235, 27] or introducing progressively larger image/feature sizes for supervision [10, 36, 152], which can also improve efficiency [150, 43]. In generative tasks, progressively selecting the camera pose is also an easy-to-difficult training strategy, optimizing from positions close to the initial viewpoint to those further away [97, 130].
Block Optimization Strategy: This strategy is often used in large-scale scene reconstruction to improve efficiency and alleviate catastrophic forgetting [154, 164, 163]. However, such paradigms are often influenced by block partitioning and training data selection. Consequently, several studies have proposed designing Primitives and Data Division strategies to mitigate workload imbalances caused by numerous empty blocks, while enhancing detail reconstruction capabilities [251]. To improve efficiency, introducing Level of Detail and hierarchical reconstruction prove effective, especially in large-scale scene processing [251, 252]. It can also achieve reconstruction by partitioning the scene into static backgrounds and dynamic objects [146, 147, 148, 175]. Additionally, this approach is applied in AIGC and Semantic Understanding, where refining submap reconstruction quality enhances overall performance [105, 227]. Unlike submaps divided by spatial regions, Gaussians can be categorized into different generations during their densification process, allowing for the application of distinct regularization strategies to each generation, effectively regulating their fluidity [201]. Categorizing Gaussians into those on smooth surfaces and independent points is also feasible for geometric representation. By designing distinct initialization and densification strategies, better representation can be achieved [242]. Additionally, some works design keyframe (or window) selection strategies based on inter-frame covisibility or geometric overlap ratio in temporal data for reconstructions [160, 250, 151, 238, 153, 172].
Robust Optimization Strategy: Introducing noise perturbations is a common method to enhance the robustness of training [174, 115, 64, 49]. Such perturbations can target camera poses, timesteps, and images, and can be regarded as a form of data augmentation to prevent overfitting. Additionally, some strategies mitigate catastrophic forgetting by avoiding continuous training from a single viewpoint [154, 156].
Distillation-based Strategy: To compress model parameters, some distillation strategies use the original 3DGS as the teacher model and a low-dimensional SH 3DGS as the student model, introducing more pseudo views to enhance the performance of the low-dimensional SH [15].
@@ VI-F Adaptive Control
VI-F1 Densification
Densification is crucial, especially for detail reconstruction. we will analyze it from the perspectives of "Where to densify" and "How to densify". Additionally, we will discuss how to avoid excessive densification.
Where to Densification: Densification techniques focus on identifying positions requiring densification, governed by gradients in the original 3DGS and extendable to dynamic scene reconstruction [170]. Regions with low opacity, silhouette, or high depth-rendered error, considered unreliable, guide densification to fill holes or improve 3D inconsistencies [250, 151, 164, 185, 26, 197]. Some approaches improve based on gradients by weighting the number of pixels covered by each Gaussian in different views to dynamically average view gradients, enhancing point cloud growth conditions [253]. Additionally, SDF value, motion masks and neighbor distance are important criteria, with locations closer to the surface, motion regions and lower compactness being more prone to densification [195, 192, 100, 179].
How to Densification: Numerous works have improved densification methods. Graph structures explore Gaussian relationships and define new Gaussians at edge centers based on proximity scores, mitigating sparse viewpoint impacts [57]. To prevent excessive Gaussian growth, the Candidate Pool Strategy stores pruned Gaussians for densification [118]. Additionally, work [254] introduces three conservation rules for visual consistency and employs integral tensor equations to model densification.